OpenStudy (anonymous):

Which of the following sets could be the sides of a right triangle? A. {2, 3,√10 } B. {3, 5, √34} C. {5, 8, 12}

OpenStudy (aum):

If a^2 + b^2 = c^2 then it is a right triangle.

OpenStudy (aum):

Square the two smaller numbers. Add them and see if it is the square of the largest number.

OpenStudy (jdoe0001):

well.. is a right-triangle that means that the LONGEST side of all three, will be equals to the SUM of the SQUARE of the other two sides

OpenStudy (jdoe0001):

so let's check the 1st one then \(\bf \sqrt{2^2+3^2}=\sqrt{10}\quad ?\) well \(\bf \sqrt{2^2+3^2}=\sqrt{10}\to \sqrt{4+9}=\sqrt{10}\to \sqrt{13}\ne \sqrt{10}\) so, it's not that one see if the others add up

OpenStudy (anonymous):

is it B?

OpenStudy (jdoe0001):

is it? well... check their square of the sides.. notice the 3rd side given is the LONGEST \(\bf \{3, 5, \sqrt{34}\}\implies \sqrt{3^2+5^2}=\sqrt{34}\quad ?\)